A quality-conscious disk manufacturer wishes to know the fraction of disks his company makes which are defective. Step 1 of 2: Suppose a sample of 7002 floppy disks is drawn. Of these disks, 6232 were not defective. Using the data, estimate the proportion of disks which are defective. Enter your answer as a fraction or a decimal number rounded to three decimal places. Step 2 of 2: Suppose a sample of 7002 floppy disks is drawn. Of these disks, 6232 were not defective. Using the data, construct the 99% confidence interval for the population proportion of disks which are defective. Round your answers to three decimal places.
Answer #1

Step 1 :

Sample Size : n=7002

Number of non-defective disks =6232

Number of defective disks : x = 7002-6232=770

Sample proportion \widehat{p} = x/n = 770/ 7002 = 0.110

Estimated proportion of disks which are defective : Sample propotion : \widehat{p} = 0.110

Step 2 :

Confidence interval for population proportion :

=\widehat{p}\pm z_{\alpha /2}\sqrt{(\widehat{p}(1-\widehat{p}))/n}

\alpha = (100-99)/100 = 0.01; \alpha/2 = 0.005

99% Confidence interval =

=\widehat{p}\pm z_{0.005}\sqrt{(\widehat{p}(1-\widehat{p}))/n}

from standard normal tables z0.005 = 2.58

99% Confidence interval for population proportion of disks which ar defective=

=\widehat{p}\pm z_{0.005}\sqrt{(\widehat{p}(1-\widehat{p}))/n}

=0.11\pm 2.58\sqrt{(0.11(1-0.11))/7002}

=0.11\pm 2.58\sqrt{(0.110\times 0.89)/7002}=0.11\pm 2.58\sqrt{\frac{0.0979}{7002}}

=0.11\pm 2.58\times 0.0037 = 0.11\pm 0.0096 =(0.1004, 0.1196)

99% Confidence interval for population proportion of disks which ar defective= (0.1004, 0.1196)